OpenStudy (anonymous):

5-20x ----- 4x-1 A) 20x - 5 B) 5 C) 4x - 1 D) -5 < Is this correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are u supposed to be dividing here?

OpenStudy (anonymous):

5-20x = 5*(1-4x) = -5*(4x-1) So (5-20x)/(4x-1) = -5*(4x-1)/(4x-1) = -5 its correct